OpenStudy (anonymous):

how do we take the partial derivatives of 3 variables.Meaning in terms of fu, fv, fw? f(u,v,w) =(w^u)(v^3)(e^(9w-(3v^2)))+5uv

OpenStudy (anonymous):

take all other variables as constants and do normal differentiation

OpenStudy (anonymous):

u get it?

OpenStudy (anonymous):

umm.. lemme check

OpenStudy (anonymous):

for df/du treat v and w as normal constants and differentiate

OpenStudy (anonymous):

fu= (w^u)*ln(w)*v^3*e^(9w-3v^2) +5v fv= (w^u)*3v^2(e^9w-3v^2) + (w^u)(v^3)(e^(9w-(3v^2) * ( 9w-6v) +5u so in fv, at the end do we take ( 9w-6v) +5u or just (-6v)+5u man it took long to type :(